### v1

@@ 3.2.2 Studied Models
With respect to the ViT, we employed the Swin Transformer (SwinT) (Liu et al., 2021) from the ViT family for comparison. Specifically, SwinT is a widely accepted ViT that serves as the backbone for various computer vision tasks. Compared to the basic ViT (Dosovitskiy et al., 2020), SwinT proposes the shifted windowing scheme to significantly reduce the computational workload of self-attention without compromising the performance. The efficiency and capability of SwinT enable us to accelerate the prototyping.
For the CNN baselines, we utilised three strong CNN models that often lead the score board on public image classification tasks, such as CIFAR (Krizhevsky et al., 2009) and ImageNet (Deng et al., 2009). Particularly, they are 1): Wide Residual Network (WRN) (Zagoruyko and Komodakis, 2016), which increases the number of filters to improve the performance of the popular ResNet (He et al., 2016); 2) EfficientNetV2 (Tan and Le, 2021), developed by neural architecture search (NAS) and offering better speed–accuracy trade-offs on ImageNet than both larger CNNs and the basic ViT; and 3) ResNeXt (Xie et al., 2017), which is specialised for image classification and outperforms the ResNet counterpart on a variety of image classification tasks.
@@
We followed the widely-accepted pretraining practice to avoid training the model from scratch. All of the chosen models are pretrained on the ImageNet dataset (Deng et al., 2009) and open-weighted. Our pretrained CNNs and SwinT are downloaded from the TorchVision library and the model repository of Hugging Face respectively.
@@ 3.3 Training Methods
Figure 4: Comparison of Classical Supervised Learning and Supervised Contrastive Learning. Top (Classical Supervised Learning). An image encoder and classifier are trained end-to-end by maximising the predicted probability (prob) of the ground-truth label. Because embeddings from different classes remain entangled, the classifier struggles to carve out a decision boundary in a poorly separated embedding space, which can hinder performance. Bottom (Supervised Contrastive Learning). First, the encoder is pretrained with a contrastive loss that pulls together same-class embeddings and pushes apart different-class embeddings. This step learns an encoder that generates well-clustered representations. A classifier is then fine-tuned on these fixed embeddings using classic supervised learning, enabling it to learn a much clearer decision boundary.
@@
To encourage more separable embeddings, we use contrastive learning (Wu et al., 2018) for a comparison with the conventional supervised learning. Built on the prior knowledge that a model should learn invariant representations for similar samples, contrastive learning pulls together any given training image, denoted as an anchor and its augmented views (positive samples), while pushing apart all other different images (negative samples) in the embedding space. Contrastive learning enables the model to learn meaningful representations for images without references to ground-truth labels.
When labels are available, contrastive learning can be more flexible for learning representations. Particularly, Supervised Contrastive Learning (SupCon) (Khosla et al., 2020) integrates the label information into the contrastive loss to force the model to learn close embeddings for instances belonging to the same class, whereas embeddings of different classes are pushed away. As such, SupCon explicitly decreases intra-class variance and boosts inter-class variance, creating tight clusters for each habitat type and clear margins between them. With more separable embeddings, our hypothesis is that SupCon can alleviate the difficulties of learning an effective classifier to discern visually similar habitats.
@@ 3.4.1 Classification Performance
[…]
@@
[…]
@@ 3.4.2 Model Interpretability
GradCAM (Selvaraju et al., 2017) is utilised for explaining model predictions. GradCAM is a widely accepted technique to generate visual explanations for predictions made by deep neural network on vision tasks. For image classification tasks, GradCAM produces a class activation map (CAM) using the gradients of the target class to highlight the crucial regions in the image for the model to predict the class. The heated regions shown in warm colours on the map suggests the class is determined by the visual features within these areas, and vice versa. GradCAM can be used to visually examine if model predictions are close to human perceptions. For instance, the GradCAM of a grassland image is expected to highlight regions where are largely occupied by grasses rather than the field surveyor.
@@ 3.5 Implementation Details
Our implementation was based on the PyTorch library. The AdamW (Loshchilov and Hutter, 2017) optimiser was chosen for model training. The learning rate, weight decay, and the batch size were tuned to $5\times 10^{-6}$, 0.05, and 16 respectively. With respect to the data preprocessing pipeline during training, images were first resized to 384 $\times$ 384 and then applied with random crop and rotation transformation before being fed to the model. The model was trained for 50 epochs as the learning performance was saturated by this point. In terms of the hardware, all of our experiments were run on a single NVIDIA 4090 GPU with 24GB memory.
For hyperparameters unique to SupCon, we followed its original settings in proposed Khosla et al. (2020). During the pretraining stage, the projection head used was a two-layer MLP with the ReLu activation. The batch size was doubled by independently applying data transformations of random crop and rotation for two times. The temperature in the contrastive loss was set to 0.1. The pretraining stopped at 50 epochs as we found the classifier performed best during the transfer learning stage with the resulting image encoder.
4 Results and Discussion
@@ 4.2 Prediction Interpretability
We attribute the illustrated interpretability of SwinT-B for habitat classification to its attention mechanism. Take the grassland as an example, SwinT learns to connect small grass regions shown across the entire image as they are visually similar to each other, and leverage the learned embeddings of connected grass regions to activate its classification layer for the prediction. Conversely, the prediction of CNN is based on confined local areas, without a learning mechanism to connect them over the entire image, resulting in its subjectiveness to local visual features, which makes CNN a poor fit for habitat classification. Combining the observations in classification performance and GradCAMs, we conclude that the capability of ViT for learning image representation encoding long-range regional dependencies and the overall contextual information of the scene, is crucial for improving the model for habitat recognition, and results in superior classification performance over CNN models.
@@ 4.3 Comparing SupCon with Supervised Learning
Table 3: Classification performance of Supervised Learning (Sup) and Supervised Contrastive Learning (SupCon) for habitat classification, based on the SwinT-B model.
Method Top-1 Acc. (%) Top-3 Acc. (%) $F_{1}$ (weighted) MCC
Sup 70.03 90.70 0.690 0.659
SupCon 71.53 (+1.50) 90.92 (+0.22) 0.706 (+0.016) 0.677 (+0.018)
@@
We applied SupCon to train the SwinT-B model to validate our hypothesis that SupCon can mitigate the misclassifications among visually similar habitats. With its capability of separating embeddings from different classes in the latent space, SupCon is anticipated to enable the model to learn a classifier that is more effective at discerning visually similar habitats compared to the classical supervised learning approach. Table 3 reports the results of our experiments to train the SwinT-B with SupCon, comparing the classification performance of SupCon with the supervised learning. SupCon outperforms its counterpart on all four classification metrics, +1.5%, +0.22%, +0.016, and +0.018 for Top-1 accuracy, Top-3 accuracy, $F_{1}$ score, MCC, respectively.
Though only modest gains are achieved on the overall classification performance, SupCon is observed to mitigate the misclassifications between challenging habitats. Specifically, Figure 7 is a delta confusion matrix that compares the normalised confusion matrices generated by SupCon and supervised learning on major habitats. With SupCon, Improved Grassland increases its correct classifications by 6% and reduces the misclassifications on Neutral Grassland by 6%. Similarly, 5% more Neutral Grassland instances are correctly predicted and its misclassifications to Improved Grassland are reduced by 2% with SupCon. In terms of FMS, SupCon significantly improves the proportion of correct classifications by 14% and decrease its misclassifications to Bog and Neutral Grassland by 5% and 6%, respectively. The correct predictions of Bracken are improved by 8% with SupCon, owing to a 12% decrease of misclassifications to Acid Grassland. Conversely, SupCon sacrifices classification performance of Dwarf Shrub Heath, decreasing its correct predictions by 4%, and fails to increase the correct classifications for Calcareous Grassland. This could be due to the fact that the Calcareous Grassland is a minor L3 habitat with a very limited number of samples in CS dataset for training the model, as shown in Figure 2). We provide more insights about the performance of SupCon with the embedding analysis in Appendix A.
@@ 4.4 Comparing the Model with Human Experts
Table 4: Performance comparison between SwinT-B and human experts on a subset drawn from the test set. Bold values are the best performance, underlined values are the best performance achieved by human experts, and performance differences between SwinT-B and the best human performance are indicated in the brackets.
Metric SwinT-B (SupCon) Expert 1 Expert 2 Expert 3
Top-1 Acc. (%) 63.92 (+0.23) 55.70 51.27 63.69
Top-3 Acc. (%) 87.34 (+7.72) 73.42 74.68 79.62
$F_{1}$ (weighted) 0.603 (-0.025) 0.541 0.529 0.628
MCC vs Ground Truth 0.599 (-0.005) 0.517 0.479 0.604
MCC vs Expert 1 0.529
MCC vs Expert 2 0.490 0.515
MCC vs Expert 3 0.546 0.511 0.553
@@
Table 4 compares the performance of our best model, SwinT-B trained with SupCon, on a subset of the test data with photographic interpretation of L3 habitats by three human experts, as specified in Section 3.4.3. Compared to the best expert (Expert 3), SwinT-B: 1) outperforms their Top-1 accuracy by 0.23% and their Top-3 accuracy significantly by 7.72%; 2) slightly underperforms on weighted $F_{1}$ and MCC by 0.025 and 0.005, respectively, primarily because it struggles to generalise to rare habitats (e.g. Montane) owing to limited training samples (see Appendix B for detailed discussion); 3) has a strong agreement with the best expert, achieving an MCC of 0.546 against them. We offer three key insights for the observation as follows:
  • Human-level performance. The developed model performs comparably with human experts on L3 habitat interpretation of field photos, demonstrating that our ViT + SupCon approach is an effective framework for automatic ground-level habitat classification. It should be noted that Expert 3 is more familiar with CS imagery and protocols, which may have given them an advantage over Experts 1 and 2.
@@
  • Inherent challenge of ground-level imagery. With a Top-1 accuracy of only ~63%, our study reveals that ground-level classification of L3 habitats from images remains difficult for both AI models and human experts. Remote sensing methods have shown better performance in multi-habitat (albeit narrower) contexts, e.g. forests (Praticò et al., 2021), wetlands (Jamali et al., 2023), grasslands (Díaz-Ireland et al., 2024). Unlike bird’s-eye remote-sensing data, ground-level images require models to understand scene depth (disentangling background from foreground), to filter out distracting elements present in the scene, and to detect often-subtle compositional differences among visually similar subtypes. These added complexities make ground-level habitat classification very challenging. Nevertheless, ground-level habitat classification remains a valuable complement to remote sensing systems since ground-level photos are easier for non-specialists to collect and use (Cao et al., 2023; Mall et al., ; Morueta-Holme et al., 2024).
@@
  • Pathways to improving ground-level habitat classification. a) The CS photos were originally taken for plot relocation, not habitat identification. Establishing clear guidelines (e.g. framing, lighting, seasonal timing) for habitat-focused imagery should enhance the training data and reveal the true potential of AI for ground-level habitat classification. b) Human surveyors use a rich collection of information and often quantitative data–information absent from 2D images–to determine habitats. For example, they take into consideration the geographical and climatic information, management context, soil type, species, etc. when conducting habitat assessment in the field. Additionally, they may visit the site during different seasons in order to record species growing at different times of the year. Incorporating multi-view or time-series photos, or auxiliary metadata (e.g. date, location), could narrow the remaining performance gap. c) Since habitat is often defined by characteristic species, extending our approach to recognise key flora (e.g. dominant shrubs, indicator plants) and integrating those outputs into the habitat classifier may also boost accuracy.

### v2

@@ 2.2.2 Studied Models
We selected representative CNN and ViT models that balance performance and efficiency for the architecture comparison. Using a size sweep to cover a range of capacities, we compared 9 models in total. With respect to the ViT architecture, we employed the Swin Transformer (SwinT) (liu2021swin) for its computational efficiency. Compared to the basic ViT (dosovitskiy2020image), SwinT proposes the shifted window scheme to significantly reduce the computational workload of self-attention, while achieving classification performance on par with or superior to basic ViTs. SwinT has become a standard representative ViT baseline for benchmarking in a number of state-of-the-art computer vision studies (liu2022convnet; zhu2023biformer; wang2023internimage). The efficiency and capability of SwinT enable us to accelerate the prototyping.
For the CNN baselines, we utilised three strong CNN models with complementary design principles, each of which has consistently ranked among the top performers on public image classification benchmarks such as CIFAR (krizhevsky2009learning) and ImageNet (deng2009imagenet). They are 1): Wide Residual Network (WRN) (zagoruyko2016wide), which increases the number of filters to improve the performance of the popular ResNet (he2016deep); 2) EfficientNetV2 (tan2021efficientnetv2), developed by neural architecture search (NAS) and offering better speed–accuracy trade-offs on ImageNet than both larger CNNs and the basic ViT; and 3) ResNeXt (xie2017aggregated) introduces cardinality—the number of parallel paths in each block—as a new dimension for scaling CNNs, and achieves superior performance to ResNet on image classification benchmarks.
@@
We followed the widely-accepted pretraining practice to avoid training the model from scratch. All of the chosen models are pretrained on the ImageNet dataset (deng2009imagenet) and open-weighted. Our pretrained CNNs and SwinT are downloaded from the TorchVision library and the model repository of Hugging Face respectively.
@@ 2.3 Training Methods
Figure 4: Comparison of Classical Supervised Learning and Supervised Contrastive Learning. Top (Classical Supervised Learning). An image encoder and classifier are trained end-to-end with cross-entropy. Because embeddings from different classes remain entangled, the classifier struggles to carve out a decision boundary in a poorly separated embedding space, which can hinder performance. Bottom (Supervised Contrastive Learning). First, the encoder is pretrained with a contrastive loss that pulls together same-class embeddings and pushes apart different-class embeddings. This step learns an encoder that generates well-clustered representations. A classifier is then fine-tuned on these fixed embeddings using classic supervised learning, enabling it to learn a much clearer decision boundary.
@@
To encourage more separable embeddings, we use contrastive learning (wu2018unsupervised) for a comparison with the conventional supervised learning. Built on the prior knowledge that a model should learn invariant representations for similar samples, contrastive learning pulls together any given training image, denoted as an anchor and its augmented views (positive samples), while pushing apart all other different images (negative samples) in the embedding space. Contrastive learning enables the model to learn meaningful representations for images without references to ground-truth labels.
When labels are available, contrastive learning can be more flexible for learning representations. Particularly, Supervised Contrastive Learning (SupCon) (khosla2020supervised) integrates the label information into the contrastive loss to force the model to learn close embeddings for instances belonging to the same class, whereas embeddings of different classes are pushed away. As such, SupCon explicitly decreases intra-class variance and boosts inter-class variance, creating tight clusters for each habitat type and clear margins between them. With more separable embeddings, our hypothesis is that SupCon can alleviate the difficulties of learning an effective classifier to discern visually similar habitats.
@@ 2.4.1 Classification Performance
$\mathit{Accuracy@1}=\frac{c^{(1)}}{s}$ (2)
$\mathit{Accuracy@3}=\frac{c^{(3)}}{s}$ (3)
$\mathit{MCC}=\frac{c^{(1)}\,s\;-\;\sum_{k}t_{k}\,p_{k}}{\sqrt{s^{2}-\sum_{k}p_{k}^{2}}\;\sqrt{s^{2}-\sum_{k}t_{k}^{2}}}$ (4)
@@
$\mathit{Precision}=\frac{\textup{TP}}{\textup{TP}+\textup{FP}}$ (5)
$\mathit{Recall}=\frac{\textup{TP}}{\textup{TP}+\textup{FN}}$ (6)
@@ 2.4.2 Model Interpretability
GradCAM (selvaraju2017grad) is utilised for explaining model predictions. GradCAM is a widely accepted technique to generate visual explanations for predictions made by deep neural network on vision tasks. For image classification tasks, GradCAM produces a class activation map (CAM) using the gradients of the target class to highlight the crucial regions in the image for the model to predict the class. The heated regions shown in warm colours on the map suggests the class is determined by the visual features within these areas, and vice versa. GradCAM can be used to visually examine if model predictions are close to human perceptions. For instance, the GradCAM of a grassland image is expected to highlight regions where are largely occupied by grasses rather than the field surveyor.
@@ 2.5 Implementation Details
Our implementation was based on the PyTorch library. The AdamW (loshchilov2017decoupled) optimiser was chosen for model training. The learning rate, weight decay, and the batch size were tuned to $5\times 10^{-6}$, 0.05, and 16 respectively. With respect to the data preprocessing pipeline during training, images were first resized to 384 $\times$ 384 and then applied with random crop and rotation transformation before being fed to the model. The model was trained for 50 epochs as the learning performance was saturated by this point. In terms of the hardware, all of our experiments were run on a single NVIDIA 4090 GPU with 24GB memory.
For hyperparameters unique to SupCon, we followed its original settings in proposed khosla2020supervised. During the pretraining stage, the projection head used was a two-layer MLP with the ReLu activation. The batch size was doubled by independently applying data transformations of random crop and rotation for two times. The temperature in the contrastive loss was set to 0.1. Contrastive pretraining was conducted for 100 epochs. However, we found the transfer performance using the pretrained image encoder plateaued after 50 epochs. Accordingly, the image encoder pretrained for 50 epochs was adopted for the transfer learning experiments.
3 Results
@@ 3.2 Prediction Interpretability
We attribute the illustrated interpretability of SwinT-B for habitat classification to its attention mechanism. Take the grassland as an example, SwinT learns to connect small grass regions shown across the entire image as they are visually similar to each other, and leverage the learned embeddings of connected grass regions to activate its classification layer for the prediction. Conversely, the prediction of CNN is based on confined local areas, without a learning mechanism to connect them over the entire image, resulting in its subjectiveness to local visual features, which makes CNN a poor fit for habitat classification. Combining the observations in classification performance and GradCAMs, we conclude that the capability of ViT for learning image representation encoding long-range regional dependencies and the overall contextual information of the scene, is crucial for improving the model for habitat recognition, and results in superior classification performance over CNN models. While SwinT generalises over a broad scene context, this does not necessarily lead to correct predictions. As shown in Figure 5 (rightmost column), SwinT misclassifies a Neutral Grassland photo to Improved Grassland. The error arises from the difficulty of distinguishing habitats with highly similar visual context. Neutral and Improved Grasslands are often subtly different and difficult to interpret in the field. Hence, addressing confusions caused by similar visual context is essential to further improve overall accuracy.
@@ 3.3 Comparing SupCon with Supervised Learning
Table 3: Performance of supervised learning (Sup) vs. supervised contrastive learning (SupCon) across architectures. Numbers are reported on the CS habitat test set. For each model, SupCon cells show the absolute gain over Sup in parentheses.
Model Top-1 Acc. (%) Top-3 Acc. (%) $\mathbf{F_{1}}$ (weighted) MCC
WRN-50-2 64.23  ( -1.36) $\downarrow$ 89.13  (+1.50) $\uparrow$ 0.627  ( -0.013) $\downarrow$ 0.593  ( -0.015) $\downarrow$
EfficientNetV2-L 67.02  ( +1.07) $\uparrow$ 90.70  ( +2.36) $\uparrow$ 0.658  ( +0.016) $\uparrow$ 0.625  ( +0.012) $\uparrow$
ResNeXt-50 64.23  ( -1.94) $\downarrow$ 88.20  ( +1.29) $\uparrow$ 0.610  ( -0.032) $\downarrow$ 0.593  ( -0.023) $\downarrow$
SwinT-B 71.53  ( +1.50) $\uparrow$ 90.92  ( +0.22) $\uparrow$ 0.706  ( +0.016) $\uparrow$ 0.677  ( +0.018) $\uparrow$
@@
[…]
For our best model, SwinT-B, SupCon is observed to mitigate the misclassifications between challenging habitats. Specifically, Figure 7 is a delta confusion matrix that compares the normalised confusion matrices generated by SupCon and supervised learning on major habitats. With SupCon, Improved Grassland increases its correct classifications by 6% and reduces the misclassifications on Neutral Grassland by 6%. Similarly, 5% more Neutral Grassland instances are correctly predicted and its misclassifications to Improved Grassland are reduced by 2% with SupCon. In terms of FMS, SupCon improves the proportion of correct classifications by 14% and decrease its misclassifications to Bog and Neutral Grassland by 5% and 6%, respectively. The correct predictions of Bracken are improved by 8% with SupCon, owing to a 12% decrease of misclassifications to Acid Grassland. Conversely, SupCon sacrifices classification performance of Dwarf Shrub Heath, decreasing its correct predictions by 4%, and fails to increase the correct classifications for Calcareous Grassland. This could be due to the fact that the Calcareous Grassland is a minor L3 habitat with a very limited number of samples in CS dataset for training the model, as shown in Figure 2). We provide more insights about the performance of SupCon with the embedding analysis in Appendix A.
For our CNN backbones, we observe that SupCon can shift the decision boundary for them rather than enlarge the inter-class margin: for Neutral vs Improved Grassland, SupCon reduces Neutral to Improved confusions but increases Improved to Neutral, yielding higher recall for Neutral at the expense of precision. A plausible explanation is the strong local prior of CNNs, which may limit the visual information expressed in the embeddings, making contrastive learning less effective to separate classes that require broader contextual cues. In contrast, SwinT-B with SupCon reduces errors in both directions, suggesting that global contextual cues leveraged by the SwinT-B are better suited to contrastive learning for the habitat classification task. Further details on the performance of CNNs with SupCon are provided in the Appendix B.
@@ 3.4 Comparing the Model with Human Experts
Table 4: Performance comparison between our best models, EfficientNetV2-L and SwinT-B, and human experts on a subset drawn from the test set. Bold values are the best performance, underlined values are the best performance achieved by human experts, and performance differences between the model and the best human performance are indicated in the brackets.
Metric EfficientNetV2-L (SupCon) SwinT-B (SupCon) Expert 1 Expert 2 Expert 3
Top-1 Acc. (%) 56.33 (-7.36) 63.92 (+0.23) 55.70 51.27 63.69
Top-3 Acc. (%) 84.18 (+4.56) 87.34 (+7.72) 73.42 74.68 79.62
$F_{1}$ (weighted) 0.535 (-0.093) 0.603 (-0.025) 0.541 0.529 0.628
MCC vs Ground Truth 0.516 (-0.088) 0.599 (-0.005) 0.517 0.479 0.604
MCC vs Expert 1 0.529 0.529
MCC vs Expert 2 0.460 0.490 0.515
MCC vs Expert 3 0.530 0.546 0.511 0.553
@@
Table 4 compares the performance of our best performed CNN and ViT trained with SupCon as shown in Table 3, on a subset of the test data with photographic interpretation of L3 habitats by three human experts, as specified in Section 2.4.3. The notably higher performance of Expert 3 likely reflects their extensive familiarity with the test data, gained through years of participation in the UK Countryside Survey. Compared to the best expert (Expert 3), SwinT-B: 1) outperforms their Top-1 accuracy by 0.23% and their Top-3 accuracy significantly by 7.72%; 2) slightly underperforms on weighted $F_{1}$ and MCC by 0.025 and 0.005, respectively, primarily because it struggles to generalise to rare habitats (e.g. Montane) owing to limited training samples (see Appendix C for detailed discussion); 3) has a strong agreement with the best expert, achieving an MCC of 0.546 against them. In contrast, our best CNN with SupCon (EfficientNetV2-L) underperforms the human experts on all headline metrics and is well below SwinT-B, indicating that studied CNNs do not yet reach expert-level performance on this task.
4 Discussion
Our experiment results show that SwinT surpasses state-of-the-art CNNs, including Wide ResNet, EfficientNetV2, and ResNext, for ground-level habitat classification. SupCon training can further boosts SwinT, achieving performance comparable to human expert. Meanwhile, the study also exposes challenges in habitat classification with ground-level imagery. We provide the following insights:
@@ 4 Discussion
  • Inherent challenge of ground-level imagery. With a Top-1 accuracy of only ~63% in human benchmark, our study reveals that ground-level classification of L3 habitats from images remains difficult for both AI models and human experts. Remote sensing methods have shown better performance in multi‐habitat (albeit narrower) contexts, e.g. forests (pratico2021machine), wetlands (jamali2023wetmapformer), grasslands (diaz2024classification). Unlike bird’s-eye remote-sensing data, ground-level images require models to understand scene depth (disentangling background from foreground), to filter out distracting elements present in the scene, and to detect often‐subtle compositional differences among visually similar subtypes. These added complexities make ground-level habitat classification very challenging. Misclassification by the AI model can trigger inappropriate actions on the ground. From an ecological perspective, misclassified habitats could lead to wrong interventions or recommendations for land management, e.g. applying mineral fertiliser to neutral/species-rich grassland. For Biodiversity Net Gain, inaccurate mapping could result in development on ecologically valuable habitats or inappropriate decisions about where to offset habitat losses following development. Therefore, addressing the challenges of ground-level image classification is critical to responsible habitat monitoring.
@@
  • Pathways to improving ground-level habitat classification. a) The CS photos were originally taken for plot relocation, not habitat identification. Establishing clear guidelines (e.g. framing, lighting, seasonal timing) for habitat‐focused imagery should enhance the training data and reveal the true potential of AI for ground‐level habitat classification. b) Human surveyors use a rich collection of information and often quantitative data–information absent from 2D images–to determine habitats. For example, they take into consideration the geographical and climatic information, management context, soil type, species, etc. when conducting habitat assessment in the field. Additionally, they may visit the site during different seasons in order to record species growing at different times of the year. Incorporating multi-view or time-series photos, or auxiliary metadata (e.g. date, location), could narrow the remaining performance gap. c) Since habitat is often defined by characteristic species, extending our approach to recognise key flora (e.g. dominant shrubs, indicator plants) and integrating those outputs into the habitat classifier may also boost accuracy.
  • Practical deployment. When deploying the model in practice, a well-designed sampling protocol is critical to ensure reliable performance. We propose: capturing clear, context-rich images with the rear camera supported by automatic quality checks; sampling multiple photo-points within each spatial unit; and acquiring several views per point (forward, nadir, and detail). Predictions are then aggregated to habitat classes with uncertainty estimates and flags for expert review, ensuring responsible use. From a user-experience perspective, network latency is the principal constraint: a reliable connection is required to transmit images and return results with minimal delay. By contrast, compute time is comparatively negligible: on our workstation (NVIDIA RTX 4090), a single SwinT-B forward pass takes milliseconds. In our web-based API prototype, end-to-end latency is typically 0.5–2.0 s per image.